OpenStudy (anonymous):

(6^3)^3 simplify the expression

OpenStudy (anonymous):

6^9

OpenStudy (anonymous):

10077696

OpenStudy (anonymous):

(a^b)^c = a^(bc)

OpenStudy (anonymous):

$$ 6^9 = 10077696 $$